NEW ANALYSIS OF THE ν_{3} AND ν_{4} BANDS OF HNO_{3} BY HIGH RESOLUTION FOURIER-TRANSFORM INFRARED SPECTRA IN THE 7.6 μm REGION: LINE POSITIONS AND LINE INTENSITIES

\underline{A. PERRIN}, J. ORPHAL, J.-M. FLAUD, S.MILJANIC, Laboratoire Inter Universitaire des Systemes Atmosphériques, CNRS, Université Paris 12, 61 Av du Général de Gaulle, 94010 Créteil Cedex France; S. KLEE, G. C. MELLAU, Justus-Liebig - Universität, Physikalisch-Chemisches Institut, Heinrich-Buff-Ring 58, D-35392 Gießen, Germany; A. GOLDMAN, Department of Physics, University of Denver, Denver, CO 80208, USA.

Nitric acid (HNO_{3}) plays an important role in the Earth’s atmosphere as a reservoir molecule of NO_{x} species. This molecule has a strong infrared signature at 7.6 μm which is one of the most commonly used for the infrared retrieval of HNO_{3} in the atmosphere together with the 11 μm one. It is therefore essential to have the best possible spectral parameters in this spectral region. The goal of this work was to get better line positions and intensities for the ν_{4} and ν_{3} interacting cold bands located at 1303.071 and 1326.182 cm^{−1} respectively as compared to those obtained few years ago ^{a}^{,}^{b}: in fact, only the total band intensity was re-sclaled during the recent update in the HITRAN database ^{c}, ^{d} HNO_{3} line list at 7.6 μm region. For this purpose a large set of new high resolution FTS spectra recorded in Giessen and in Denver were analysed. For the line positions and line intensities calculation, the theoretical model accounts for the vibration-resonances which link the energy levels of the 4^{1} and 3^{1} bright vibrational states together with those of several dark interacting states (9^{3} for example, among others). Very preliminary results will be presented.

aA. Perrin, O. Lado-Bordowski, and A.Valentin, Mol. Physics. 67, 249 (1989)

bA. Perrin, J.-M. Flaud, C.Camy-Peyret, V. Jaouen, R. Farrenq, G. Guelachvili, Q. Kou, F. Le-Roy, M..Morillon-Chapey, J.

Orphal, M. Badaoui, J.-Y. Mandin, and V.Dana, J. Mol. Spectrosc. 160, 524 (1993)

chttp://cfa-www.harvard.edu/hitran/

dJ.-M. Flaud, G. Brizzi, M. Carlotti, A. Perrin, and M. Ridolfi, Atmos.Chem.Phys.6, 5037-5048 (2006)